OpenStudy (anonymous):

How do I find out if the square root of a number is irrational. I need this step by step so I can understand it! The numbers I want to find out this for is 81, 64, 16, and 10. @Hero Can you help, please? I don't want someone to give me the answer, I really wanna know this!

OpenStudy (anonymous):

@mathstudent55 @jim_thompson5910 @ganeshie8 @Compassionate @johnweldon1993

hero (hero):

I'll explain it a different way.

OpenStudy (anonymous):

Please!

hero (hero):

You still have to understand both ways of doing it either way. But there's a way to explain it with fractions.

OpenStudy (anonymous):

ok

hero (hero):

First of all, this is a rule (something to memorize and not necessarily to understand just yet): \(\sqrt{n} = n^{1/2}\)

OpenStudy (anonymous):

ok..i get that..

hero (hero):

\(\sqrt{n^2} = (n^2)^{1/2} = n^{(2*1/2)} = n^{2/2} = n^1 = n\)

hero (hero):

In short, the square root of n squared is n: \(\sqrt{n^2} = n\)

OpenStudy (anonymous):

So.... √81^2?

hero (hero):

We're not there yet

hero (hero):

You want to get to the goal line but you haven't ran the marathon yet.

OpenStudy (anonymous):

k..

hero (hero):

Here's what you (should) know: \(1 \times 1 = 1\) \(2 \times 2 = 4\) \(3 \times 3 = 9\) \(4 \times 4 = 16\) \(5 \times 5 = 25\) \(6 \times 6 = 36\) \(7 \times 7 = 49\) \(8 \times 8 = 64\) \(9 \times 9 = 81\)

hero (hero):

Similarly: \(1^2 = 1\) \(2^2 = 4\) \(3^2 = 9\) \(4^2 = 16\) \(5^2 = 25\) \(6^2 = 36\) \(7^2 = 49\) \(8^2 = 64\) \(9^2 = 81\)

hero (hero):

Are you with me so far?

OpenStudy (anonymous):

Yes! :)

hero (hero):

Also: \(\dfrac{1^2}{1} = 1\) \(\dfrac{2^2}{2} = 2\) \(\dfrac{3^2}{3} = 3\) \(\dfrac{4^2}{4} = 4\) \(\dfrac{5^2}{5} = 5\) \(\dfrac{6^2}{6} = 6\) \(\dfrac{7^2}{7} = 7\) \(\dfrac{8^2}{8} = 8\) \(\dfrac{9^2}{9} = 9\) Again, what you should know. If not, commit to memory.

OpenStudy (anonymous):

kk...

OpenStudy (anonymous):

81 is rational?

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

81 - Rational 64 - Rational 16 - Rational 10 - Irrational Is that right? @Hero

OpenStudy (anonymous):

@Hero

hero (hero):

All of them are rational.

hero (hero):

20/2 = 10

hero (hero):

Every integer is rational

OpenStudy (anonymous):

...All of them can't be rational... Which of the following is an irrational number? A. √81 B. √64 C. √16 D. √10 @Hero

hero (hero):

That's not what you wrote the first time

hero (hero):

\(10 \ne \sqrt{10}\)

OpenStudy (anonymous):

That is what I wrote. I wanted to find out how to find out which one was irrational. -_-

hero (hero):

10 is rational. \(\sqrt{10}\) is irrational. You're confusing yourself.

hero (hero):

I know exactly what you wrote.

OpenStudy (anonymous):

Woops, sorry. I admit defeat. :) I understand now. Sorry! Thanks for calling me out on it. :)

OpenStudy (anonymous):

Thanks @hero Your the best!! :)

hero (hero):

You know \(\sqrt{2}\) is irrational. Well \(\sqrt{10} = \sqrt{5 \times 2} = \sqrt{5}\sqrt{2}\) \(\sqrt{10}\) includes \(\sqrt{2}\)

OpenStudy (anonymous):

Would √11 be irrational?

hero (hero):

11 is a prime number. The square root of a prime number is irrational.